OpenStudy (anonymous):

Step-by-step How do i solve (9x-27)/9x

jhonyy9 (jhonyy9):

9(x-3)/9x = (x-3)/x

jhonyy9 (jhonyy9):

x-3=0 x=3

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

because 27=9*3

jhonyy9 (jhonyy9):

than you can simplify by 9

OpenStudy (anonymous):

OK

jhonyy9 (jhonyy9):

yw